OpenStudy (anonymous):

MEDAL AND FAN: HELP PLEASE!!! Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). (Answer Choices): A: Parallel y=-x-1 Perpendicular y=x+2 B: Parallel y=x-1 Perpendicular y=-x+1 C: Parallel y=x+1 Perpendicular y=-x-3 D: Parallel y=2x-2 Perpendicular y=-2x-1 I think it may be D?

OpenStudy (michele_laino):

Hint: parallel lines have the same slope for example, here are two parallel lines: \(y=2x-3,\;\;\;\;y=2x+7\)

OpenStudy (anonymous):

I'm confused ;-;

OpenStudy (michele_laino):

whereas, here are two perpendicular lines: \(y=2x+5,\;\;\;\;\; y=(-1/2)x-9\) as we can see the product of the slopes is: \[\Large 2 \cdot \left( { - \frac{1}{2}} \right) = - 1\]

OpenStudy (michele_laino):

in general, the equation slope-intercept of a line is: \(y=mx+q\), where \(m\) is the slope

OpenStudy (michele_laino):

I think that option D is wrong

OpenStudy (anonymous):

Ah, ok. So its B? I think I understand. I'm sorry, math is my worst subject :/

OpenStudy (michele_laino):

let's consider option C, for example: there the parallel line is: \(y=x+1\), now I ask what is the slope of such line?

OpenStudy (michele_laino):

hint: 

OpenStudy (michele_laino):

and the slope of the given line \(y=x+5\) is also \(1\), so such lines: \(y=x+1\), and \(y=x+5\) are parallel lines

OpenStudy (michele_laino):

next: the slope of the line (option C) \(y=-x-3\) is \(-1\), and the product of that slope with the slope of the given line \(y=x+5\), is: \((-1) \cdot 1=-1\), so the line \(y=-x-3\) is perpendicular to the given line

OpenStudy (anonymous):

I'm sorry for the slow replies,answering other questions at the same time. I'm trying to get done with my exam before 12am haha but it's all so confusing.

OpenStudy (michele_laino):

now, the line \(y=x+1\) passes at point \((-2,-1)\), since if I replace \(x=-2\) into that equation, I get: \(y=-2+1=-1\)

OpenStudy (michele_laino):

no problem :)

OpenStudy (michele_laino):

now, if I replace \(x=-2\) into the equatiopn \(y=-x-3\), I get: \(y=-(-2)-3=...?\) please continue

OpenStudy (michele_laino):

equation*

OpenStudy (michele_laino):

hint: what is \(-(-2)-3=...?\)

OpenStudy (anonymous):

Wait so I was right?

OpenStudy (michele_laino):

no, since option D contains two lines whose slopes are not the right slopes

OpenStudy (michele_laino):

furthermore option B contain the lines \(y=x-1\) which doesn't pass at point \((-2,-1)\)

OpenStudy (michele_laino):

oops.. line*

OpenStudy (anonymous):

its c